solve the differential equation dy/dx + (x/1-x^2)y = x(y)^1/2 , y(0)=1
1
Expert's answer
2020-03-19T06:18:16-0400

This is the Bernoully equation

"y=uv,y'=u'v+uv'," then "v(u'+\frac{x}{1-x^2}u)+(uv'-xu^\frac{1}{2}v^\frac{1}{2})=0"

"\int \frac{du}{u}=-\int \frac{x}{1-x^2}dx" hence "u=(1-x^2)^\frac{1}{2}," then

"(1-x^2)^\frac{1}{2}v'-x(1-x^2)^\frac{1}{4}v^\frac{1}{2}=0" hence "\int \frac{dv}{v^\frac{1}{2}}=\int \frac{x}{(1-x^2)^\frac{1}{4}}dx" hence

"v=(-\frac{(1-x^2)^\frac{3}{4}}{3}+C)^2" then "y=(1-x^2)^\frac{1}{2}v"

as "y(0)=1" hence "C=4\/3" then "y(x)=\frac{1}{9}(1-x^2)^\frac{1}{2}(4-(1-x^2)^\frac{3}{4})^2"